# Effective and Efficient Cross-City Traffic Knowledge Transfer: A Privacy-Preserving Perspective

### v1

@@ Abstract.
Traffic prediction targets forecasting future traffic conditions using historical traffic data, serving a critical role in urban computing and transportation management. To mitigate the scarcity of traffic data while maintaining data privacy, numerous Federated Traffic Knowledge Transfer (FTT) approaches have been developed. These methods integrate transfer learning and federated learning to transfer traffic knowledge from data-rich source cities to data-scarce target cities, enhancing traffic prediction capabilities for the latter. However, current FTT approaches face challenges such as privacy leakage, cross-city data distribution discrepancies, low data quality, and inefficient knowledge transfer, limiting their privacy protection, effectiveness, robustness, and efficiency in real-world applications.
To this end, we propose FedTT, an effective, efficient, and privacy-aware cross-city traffic knowledge transfer framework that transforms the traffic data domain from the data-rich cities and trains traffic models using the transformed data for the data-scarce cities. First, to safeguard data privacy, we propose a traffic secret transmission method that securely transmits and aggregates traffic domain-transformed data from source cities using a lightweight secret aggregation approach. Second, to mitigate the impact of traffic data distribution discrepancies on model performance, we introduce a traffic domain adapter to uniformly transform traffic data from the source cities’ domains to that of the target city. Third, to improve traffic data quality, we design a traffic view imputation method to fill in and predict missing traffic data. Finally, to enhance transfer efficiency, FedTT is equipped with a federated parallel training method that enables the simultaneous training of multiple modules. Extensive experiments using 4 real-life datasets on 3 mainstream traffic prediction tasks demonstrate that the proposed FedTT framework outperforms the 14 state-of-the-art baselines. All the code and data are available at https://anonymous.4open.science/r/FedTT.
@@ 1. Introduction
In the spatio-temporal databases, numerous TP methods have been proposed (Zhao et al., 2023; Ji et al., 2023; Jiang et al., 2023), which typically rely on a large number of labeled data to train high-performing traffic models. However, urban traffic data is often insufficient or unavailable (Lin et al., 2018; Wang et al., 2019, 2022a), particularly in emerging cities, such as developing regions in Midwestern United States (us, 2021), where sensors are newly deployed or data collection is still in its early stages. In such cases, training traffic models becomes particularly challenging and prone to overfitting, limiting the accuracy of TP tasks (Jin et al., 2023; Mo and Gong, 2023). To address this challenge, Transfer Learning (TL) (Thirumuruganathan et al., 2018; Di et al., 2019) has been widely adopted to enhance the traffic modeling capabilities of the data-scarce target city by transferring traffic knowledge from data-rich source cities.
@@
To enable data-scarce target cities to acquire diverse traffic knowledge and improve the robustness of trained models, existing TL-based TP methods (Liu et al., 2023; Lu et al., 2022; Tang et al., 2022) tend to perform joint traffic knowledge transfer from multiple source cities. However, these approaches typically rely on centralized frameworks, which involve sharing and exchanging traffic data among cities without considering data privacy. These methods assume that different cities can directly share their original traffic data, which is overly idealistic in reality. This is because, the direct sharing of city traffic data risks privacy leakage (Liu et al., 2020; Meng et al., 2021; Yang et al., 2024) due to the presence of personal and sensitive information, such as license plate numbers and travel trajectories. Moreover, with growing privacy concerns, many laws and regulations, such as GDPR (gdp, 2025) and CCPA (ccp, 2025), mandate data collectors to minimize non-essential data transmission and avoid centralized storage of private information. Therefore, maintaining the decentralization of traffic data in TP is critical to safeguarding privacy. As shown in Fig. 1(a), PeMSD4 (pem, 2024), FT-AED (Coursey et al., 2024), HK-Traffic (hk, 2024), and PeMSD8 (pem, 2024) are four commonly used traffic datasets, which correspond to San Francisco, Nashville, Hong Kong, and San Bernardino, respectively. Among these, San Francisco, Nashville, and Hong Kong represent the source cities, while San Bernardino serves as the target city. Due to legal restrictions, traffic data cannot be shared or exchanged among cities, meaning each city can only access its local data. Under these constraints, transferring traffic knowledge from these three source cities to the target city without sharing and exchanging raw traffic data becomes challenging and critical.
@@
Challenge 1: How to effectively protect data privacy during federated traffic knowledge transfer? Although existing methods utilize federated learning to avoid direct data exchange, there is still a potential risk of data privacy leakage. This arises because these methods require the uploading of training gradients or model parameters for aggregation during federated training, which may allow attackers to infer raw data by inference attacks (Gao et al., 2024; Wang et al., 2024a; Zheng et al., 2024), as depicted in Fig. 2(a). To mitigate this risk, a straightforward approach is to apply privacy-preserving techniques such as Homomorphic Encryption (HE) (Rivest et al., 1978) and Differential Privacy (DP) for secure aggregation on the uploaded data during federated training. However, HE could introduce significant additional computation and communication overheads, which diminishes training efficiency, while DP lowers data utility and thus decreases model accuracy, as proved by previous studies (Wang et al., 2024b; Tawose et al., 2023; Dong et al., 2024). Therefore, ensuring the protection of traffic data privacy without compromising training efficiency and model accuracy remains a significant challenge.
@@
Challenge 4: How to reduce computation and communication overheads during federated traffic knowledge transfer? Existing federated traffic knowledge transfer approaches are inefficient due to several factors. First, their two-stage knowledge transfer process operates serially, which is inefficient. Second, to achieve good prediction accuracy, existing traffic prediction models are designed with high complexity and a large number of parameters (Zhao et al., 2023; Ji et al., 2023; Jiang et al., 2023), leading to significant computation and communication overheads, further limiting the efficiency of traffic knowledge transfer. As illustrated in Fig. 2(d), we report the running time (in hours) of the SOTA FTT methods (i.e., T-ISTGNN and pFedCTP) and a non-transfer method (i.e., Gated Recurrent Unit (GRU) (Chung et al., 2014) model is trained directly on the target city’s data without traffic knowledge transfer). Compared to the non-transfer method, the existing FTT methods exhibit considerably higher computation and communication overheads, with running times exceeding 40$\times$ that of the non-transfer method. As a result, how to reduce computation and communication overheads in traffic knowledge transfer to improve the efficiency of FTT is the fourth unaddressed challenge.
@@ 2.2. Traffic Knowledge Transfer
First, STT (Jin et al., 2022; Fang et al., 2022; Wang et al., 2019; Lin et al., 2018; Yao et al., 2023; Huang et al., 2023; Li et al., 2023; Chen et al., 2022a; Wang et al., 2022a) studies focus on transferring traffic knowledge from a single source city to a target city. Specifically, TransGTR (Jin et al., 2023) jointly learns transferable structure generators and forecasting models across cities to enhance prediction performance in data-scarce target cities. Next, CityTrans (Ouyang et al., 2024) leverages adaptive spatio-temporal knowledge and domain-invariant features for accurate traffic prediction in data-scarce cities. Additionally, MGAT (Mo and Gong, 2023) uses a meta-learning algorithm to extract multi-granular regional features from each source city to improve the effectiveness of traffic knowledge transfer. However, the performance of these STT methods can be significantly compromised when there are substantial differences in traffic data distribution between the source and target cities.
@@ 3. Problem Definitions
$X_{t}$, $\mathcal{X}$ The traffic data at time $t$ and a set of traffic data $\{X_{1},X_{2},\ldots\}$
@@
Definition 1 (Road Network). The road network is a weighted graph $\mathcal{G}=(\mathcal{M},\mathcal{E},A)$, where =$\mathcal{M}=\{m_{1},m_{2},\dots\}$ is the set of sensors, $\mathcal{E}\subseteq\mathcal{M}\times\mathcal{M}$ is the set of edges, and $A\in\mathbb{R}^{|\mathcal{M}|\times|\mathcal{M}|}$ is the weighted adjacency matrix of edges. Here, $m_{i}$ denotes the sensor with index $i$.
@@
Problem Formulation (FTT). In federated learning, multiple clients $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{n}\}$ collaboratively train a global model using their local data. In the first stage, FTT trains a traffic model $\theta_{\textit{TP}}$ to learn traffic knowledge from source cities =$\mathcal{R}=\{R_{1},R_{2},\ldots,R_{n}\}$, where each source city $R_{i}$ corresponds to a client $c_{i}$ shown below:
@@ 4.1. Traffic View Imputation
Design Motivation. Existing federated traffic transfer methods often overlook the challenges associated with low-quality traffic data, especially when missing data is prevalent, thereby significantly undermining the performance of traffic knowledge transfer models. Drawing inspiration from data imputation techniques (Yuan et al., 2024; Chen et al., 2022b; Peng et al., 2023) for addressing missing data, we propose the Traffic View Imputation (TVI) module, as shown in Fig. 4. This module enhances traffic data quality by completing and predicting missing traffic data through a comprehensive exploration of the spatial and temporal dependencies inherent in traffic data, as formally shown below:
@@
where $v_{t}^{i}[j][k]$ represents the traffic data of the $k$-th sensor in the $j$-th combination within the $i$-level traffic subview at time $t$, ${(v_{t}^{i}[j][k])}^{\top}\in\mathbb{R}^{1\times F_{1}}$ denotes the transpose matrix of $v_{t}^{i}[j][k]$, and $H(v_{t}^{i}[j][k])\in\mathbb{R}^{K\times F_{2}\times 1}$ represents the multi-head feature of the sensor corresponding to $v_{t}^{i}[j][k]$. The term $H(v_{t}^{i}[j][k])\cdot{(v_{t}^{i}[j][k])}^{\top}\in\mathbb{R}^{K\times F_{2}\times F_{1}}$ is the matrix multiplication of $H(v_{t}^{i}[j][k])$ and ${(v_{t}^{i}[j][k])}^{\top}$. Additionally, $\theta_{E}$ is a fully connected layer with $K\times F_{2}\times F_{1}$ input dimensions and ||×$|\mathcal{M}|\times F_{1}$ output dimensions.
@@ 4.2. Traffic Domain Adapter
$\mathop{\min}_{\theta_{\textit{Dis}}}\frac{1}{|\mathcal{X}^{\textit{RS}}|}\sum^{|\mathcal{X}^{\textit{RS}}|}_{t=1}\begin{cases}\begin{aligned} &-log(P(X_{t}^{\textit{RS}}\in\mathcal{X}^{R\to S})),\ \ if\ X_{t}^{\textit{RS}}\in\mathcal{X}^{R\to S}\\ &-log(P(X_{t}^{\textit{RS}}\in\mathcal{X}^{S}))\ \ \ \ \ ,\ \ if\ X_{t}^{\textit{RS}}\in\mathcal{X}^{S}\end{aligned}\end{cases}$
@@ 4.3. Traffic Secret Transmission
Design Motivation. Existing FTT works upload training gradients or model parameters for model aggregation in the federated training, where attackers derive the traffic data from the uploaded model parameters or training gradients through inference attacks (Gao et al., 2024; Wang et al., 2024a; Zheng et al., 2024). Although some techniques such as Homomorphic Encryption (HE) (Rivest et al., 1978) and Differential Privacy (DP) (Dwork et al., 2006) can be employed for federated secure aggregation, HE introduces significant computational and communication overheads, reducing training efficiency, while DP reduces the data utility, leading to lower model accuracy. In contrast, to protect traffic data privacy, we design the Traffic Secret Transmission (TST) module to securely transmit and aggregate the transformed data from source cities by a lightweight traffic secret aggregation method without sacrificing the training efficiency or model accuracy, as shown in Fig. 5 (③–④).
@@
where $\mathcal{X}^{R_{i}\to S}_{(r)}$ is the transformed data from the source city $R_{i}$ at round $r$ , and $\overline{\mathcal{X}}^{\mathcal{R}\to S}_{(r)}$ is the aggregated data at round $r$. In this way, it not only protects traffic data privacy as information for individual source cities can not be inferred or derived from the aggregated data (Wang et al., 2024b; Dong et al., 2024) but also reduces the influence of specific traffic patterns of source cities on traffic models when the target city uses the aggregated data for model training.
@@
Finally, the server gets the aggregated data at round $r$ based on the aggregated data $\overline{\mathcal{X}}_{(r-1)}^{\mathcal{R}\to S}$ at round $r-1$, as shown below:
@@
$\mathop{\min}_{\theta_{\textit{Dis}}^{R_{i}}}\frac{1}{|\mathcal{X}^{R_{i}S}|}\sum^{|\mathcal{X}^{R_{i}S}|}_{t=1}\begin{cases}\begin{aligned} &-log(P(X^{R_{i}S}_{t}\in\mathcal{X}^{R_{i}\to S})),\ \ if\ X^{R_{i}S}_{t}\in\mathcal{X}^{R_{i}\to S}\\ &-log(P(X^{R_{i}S}_{t}\in\mathcal{X}^{\mathcal{R}\to S})),\ \ if\ X^{R_{i}S}_{t}\in\mathcal{X}^{\mathcal{R}\to S}\end{aligned}\end{cases}$
@@ 4.4. Federated Parallel Training
Input: clients $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{n}\}$ (i.e., source cities $\mathcal{R}=$
$\{R_{1},R_{2},\ldots,R_{n}\}$) and the server $s$ (i.e., the target city $S$).
1 In the client $c_{i}\in\mathcal{C}$:
2     $\widetilde{D}^{R_{i}}\stackrel{{\scriptstyle\theta_{\textit{TVI}}}}{{\longleftarrow}}D^{R_{i}}$
3     for training round $r=1,2,\ldots$ do
4          $\mathcal{X}^{R_{i}\to S}_{(r)}\stackrel{{\scriptstyle\theta_{\textit{Gen}}^{R_{i}}}}{{\longleftarrow}}\widetilde{D}^{R_{i}}$
5          $\overline{\mathcal{X}}^{\mathcal{R}\to S}_{(r-1)}\leftarrow\textit{GetAggregated}(s,r)$
6          $\textit{Classify}(\theta_{\textit{Dis}}^{R_{i}},\overline{\mathcal{X}}^{\mathcal{R}\to S}_{(r-1)},\mathcal{X}^{R_{i}\to S}_{(r)})$
7          $\mathcal{X}^{(\mathcal{R}\to S,R_{i})}_{(r)}\leftarrow\overline{\mathcal{X}}^{\mathcal{R}\to S}_{(r-1)}+\mathcal{X}^{R_{i}\to S}_{(r)}-\mathcal{X}^{R_{i}\to S}_{(r-1)}$
8          $\textit{Send}(s,\mathcal{X}^{(\mathcal{R}\to S,\ R_{i})}_{(r)})$
9
10
11     In the server $s$:
12          for training round $r=1,2,\ldots$ do
13               $\{\mathcal{X}^{(\mathcal{R}\to S,\ R_{1})}_{(r)},\mathcal{X}^{(\mathcal{R}\to S,\ R_{2})}_{(r)},\ldots\}\leftarrow\textit{GetAggregated}(\mathcal{C},r)$
14               $\overline{\mathcal{X}}^{\mathcal{R}\to S}_{(r)}\leftarrow\sum^{n}_{i=1}\mathcal{X}^{(\mathcal{R}\to S,R_{i})}_{(r)}-(n-1)*\overline{\mathcal{X}}^{\mathcal{R}\to S}_{(r-1)}$
15               $\textit{Classify}(\theta_{\textit{Dis}},\overline{\mathcal{X}}^{\mathcal{R}\to S}_{(r)},\mathcal{X}^{S})$
16               $\widetilde{\mathcal{X}}^{S}_{(r)}\stackrel{{\scriptstyle\theta_{\textit{TP}}}}{{\longleftarrow}}(\overline{\mathcal{X}}^{\mathcal{R}\to S}_{(r)},\mathcal{X}^{S}),$
17               $\textit{Send}(\mathcal{C},\overline{\mathcal{X}}^{\mathcal{R}\to S}_{(r)})$
18
Algorithm 1 The training of the FedTT framework

### v4

@@ Abstract.
Traffic prediction (TP) is a core task in urban computing, aiming to forecast future traffic conditions from historical observations. To overcome the scarcity of traffic data in emerging cities, recent studies have explored Federated Traffic Knowledge Transfer (FTT), which leverages data-rich source cities to assist data-scarce target cities without raw data sharing. However, existing FTT approaches are limited by three unresolved challenges: (i) potential privacy leakage since gradients or parameters generated during federated computing can still be inverted, (ii) severe cross-city distribution discrepancies that reduce transfer effectiveness, and (iii) low data quality caused by missing or unreliable sensor readings. To address these challenges, we propose FedTT, a novel federated framework for cross-city traffic knowledge transfer with privacy-preserving. FedTT introduces three innovations: (i) a lightweight Traffic Secret Aggregation (TSA) protocol that achieves secure knowledge aggregation without sacrificing efficiency or accuracy; (ii) a Traffic Domain Adapter (TDA) that explicitly aligns heterogeneous source–target distributions for more effective transfer, and (iii) a Traffic View Imputation (TVI) method that leverages spatio-temporal dependencies to complete missing traffic data robustly. Extensive experiments on four real-world datasets show that FedTT achieves significant improvements over 18 state-of-the-art baselines, consistently reducing prediction error while maintaining strong privacy protection.
@@ 1. Introduction
In the spatio-temporal databases, numerous TP methods have been proposed (Zhao et al., 2023; Ji et al., 2023; Jiang et al., 2023), which typically rely on a large number of labeled data to train high-performing traffic models. However, urban traffic data is often insufficient or unavailable (Lin et al., 2018; Wang et al., 2019, 2022a), particularly in emerging cities, such as developing regions in Midwestern United States (37), where sensors are newly deployed or data collection is still in its early stages. In such cases, training traffic models becomes particularly challenging and prone to overfitting, limiting the accuracy of TP tasks (Jin et al., 2023; Mo and Gong, 2023). To address this challenge, Transfer Learning (TL) (Thirumuruganathan et al., 2018; Di et al., 2019) has been widely adopted to enhance the traffic modeling capabilities of the data-scarce target city by transferring traffic knowledge from data-rich source cities.
@@
To enable data-scarce target cities to acquire diverse traffic knowledge and improve the robustness of trained models, existing TL-based TP methods (Liu et al., 2023; Lu et al., 2022; Tang et al., 2022) tend to perform joint traffic knowledge transfer from multiple source cities. However, these approaches typically rely on centralized frameworks, which involve sharing and exchanging traffic data among cities without considering data privacy. These methods assume that different cities can directly share their original traffic data, which is overly idealistic in reality. This is because, the direct sharing of city traffic data risks privacy leakage (Liu et al., 2020; Meng et al., 2021; Yang et al., 2024) due to the presence of personal and sensitive information, such as license plate numbers and travel trajectories. Moreover, with growing privacy concerns, many laws and regulations, such as GDPR (15) and CCPA (1), mandate data collectors to minimize non-essential data transmission and avoid centralized storage of private information. Therefore, maintaining the decentralization of traffic data in TP is critical to safeguarding privacy. As shown in Fig. 1(a), PeMSD4 (2), FT-AED (Coursey et al., 2024), HK-Traffic (53), and PeMSD8 (2) are four commonly used traffic datasets, which correspond to San Francisco, Nashville, Hong Kong, and San Bernardino, respectively. Among these, San Francisco, Nashville, and Hong Kong represent the source cities, while San Bernardino serves as the target city. Due to legal restrictions(e.g., GDPR (15) and CCPA (1)), traffic data cannot be shared or exchanged among cities, meaning each city can only access its local data. Under these constraints, transferring traffic knowledge from these three source cities to the target city without sharing and exchanging raw traffic data becomes challenging and critical.
@@
Challenge 1: How to effectively protect data privacy during federated traffic knowledge transfer? Although existing methods utilize federated learning to avoid direct data exchange, there is still a potential risk of data privacy leakage. This arises because these methods require the uploading of training gradients or model parameters for aggregation during federated training, which may allow attackers to infer raw data by inference attacks (Gao et al., 2024; Wang et al., 2024b; Zheng et al., 2024), as depicted in Fig. 2(a). To mitigate this risk, a straightforward approach is to apply privacy-preserving techniques such as Homomorphic Encryption (HE) (Rivest et al., 1978) and Differential Privacy (DP) for secure aggregation on the uploaded data during federated training. However, HE could introduce significant additional computation and communication overheads, which diminishes training efficiency, while DP lowers data utility and thus decreases model accuracy, as proved by previous studies (Wang et al., 2024a; Tawose et al., 2023; Dong et al., 2024). Therefore, ensuring the protection of traffic data privacy without compromising training efficiency and model accuracy remains a significant challenge.
@@
[…]
@@ 2.2. Traffic Knowledge Transfer
First, STT (Jin et al., 2022; Fang et al., 2022; Wang et al., 2019; Lin et al., 2018; Yao et al., 2023; Huang et al., 2023; Li et al., 2023; Chen et al., 2022b; Wang et al., 2022a) studies focus on transferring traffic knowledge from a single source city to a target city. Specifically, TransGTR (Jin et al., 2023) jointly learns transferable structure generators and forecasting models across cities to enhance prediction performance in data-scarce target cities. Next, CityTrans (Ouyang et al., 2024) leverages adaptive spatio-temporal knowledge and domain-invariant features for accurate traffic prediction in data-scarce cities. Additionally, MGAT (Mo and Gong, 2023) uses a meta-learning algorithm to extract multi-granular regional features from each source city to improve the effectiveness of traffic knowledge transfer. However, the performance of these STT methods can be significantly compromised when there are substantial differences in traffic data distribution between the source and target cities.
@@ 3. Problem Definitions
$_{t}, 𝒳\mathcal{X}$ The traffic data at time $t$ and a set of traffic data $\{X_{1},X_{2},\ldots\}$
@@
Definition 1 (Road Network). The road network is a weighted graph $\mathcal{G}=(\mathcal{M},\mathcal{E},A)$, where $\mathcal{M}=\{m_{1},m_{2},\dots\}$ is the set of sensors, $\mathcal{E}\subseteq\mathcal{M}\times\mathcal{M}$ is the set of edges, and $A\in\mathbb{R}^{|\mathcal{M}|\times|\mathcal{M}|}$ is the weighted adjacency matrix of edges. Here, $m_{i}$ denotes the sensor with index $i$.
@@
Problem Formulation (FTT). In federated learning, multiple clients $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{n}\}$ collaboratively train a global model using their local data. In the first stage, FTT trains a traffic model $\theta_{\textit{TP}}$ to learn traffic knowledge from source cities $\mathcal{R}=\{R_{1},R_{2},\ldots,R_{n}\}$, where each source city $R_{i}$ corresponds to a client $c_{i}$ shown below:
@@ 4.1. Traffic View Imputation
Design Motivation. Existing federated traffic transfer methods often overlook the challenges associated with low-quality traffic data, especially when missing data is prevalent, thereby significantly undermining the performance of traffic knowledge transfer models. Drawing inspiration from data imputation techniques (Yuan et al., 2024; Chen et al., 2022a; Peng et al., 2023) for addressing missing data, we propose the Traffic View Imputation (TVI) module, as shown in Fig. 4. This module enhances traffic data quality by completing and predicting missing traffic data through a comprehensive exploration of the spatial and temporal dependencies inherent in traffic data, as formally shown below:
@@
where $v_{t}^{i}[j][k]$ represents the traffic data of the $k$-th sensor in the $j$-th combination within the $i$-level traffic subview at time $t$, ${(v_{t}^{i}[j][k])}^{\top}\in\mathbb{R}^{1\times F_{1}}$ denotes the transpose matrix of $v_{t}^{i}[j][k]$, and $H(v_{t}^{i}[j][k])\in\mathbb{R}^{K\times F_{2}\times 1}$ represents the multi-head feature of the sensor corresponding to $v_{t}^{i}[j][k]$. The term $H(v_{t}^{i}[j][k])\cdot{(v_{t}^{i}[j][k])}^{\top}\in\mathbb{R}^{K\times F_{2}\times F_{1}}$ is the matrix multiplication of $H(v_{t}^{i}[j][k])$ and $(v_{t}^{i}[j][k])$$^{\top}$. Additionally, $\theta_{E}$ is a fully connected layer with $K\times F_{2}\times F_{1}$ input dimensions and $|\mathcal{M}|\times F_{1}$ output dimensions.
@@ 4.2. Traffic Domain Adapter
[…]
@@ 4.3. Traffic Secret Transmission
Design Motivation. Existing FTT works upload training gradients or model parameters for model aggregation in the federated training, where attackers derive the traffic data from the uploaded model parameters or training gradients through inference attacks (Gao et al., 2024; Wang et al., 2024b; Zheng et al., 2024). Although some techniques such as Homomorphic Encryption (HE) (Rivest et al., 1978) and Differential Privacy (DP) (Dwork et al., 2006) can be employed for federated secure aggregation, HE introduces significant computational and communication overheads, reducing training efficiency, while DP reduces the data utility, leading to lower model accuracy. In contrast, to protect traffic data privacy, we design the Traffic Secret Transmission (TST) module to securely transmit and aggregate the transformed data from source cities by a lightweight traffic secret aggregation method without sacrificing the training efficiency or model accuracy, as shown in Fig. 5 (③–④).
@@
where $\mathcal{X}^{R_{i}\to S}_{(r)}$ is the transformed data from the source city $R_{i}$ at round $r$ , and $\overline{\mathcal{X}}^{\mathcal{R}\to S}_{(r)}$ is the aggregated data at round $r$. In this way, it not only protects traffic data privacy as information for individual source cities can not be inferred or derived from the aggregated data (Wang et al., 2024a; Dong et al., 2024) but also reduces the influence of specific traffic patterns of source cities on traffic models when the target city uses the aggregated data for model training.
@@
Finally, the server gets the aggregated data at round $r$ based on the aggregated data $\overline{\mathcal{X}}_{(r-1)}^{\mathcal{R}\to S}$ at round r1r-1, as shown below:
@@
[…]
@@ 4.4. Federated Parallel Training
Input: Server $s$ (target city $S$); local data $\mathcal{X}^{R_{i}}$; models $\theta_{\text{TVI}},\theta_{\text{Gen}}^{R_{i}},\theta_{\text{Dis}}^{R_{i}}$
$\widetilde{\mathcal{X}}^{R_{i}}\leftarrow\textsc{Complete}(\theta_{\text{TVI}},\mathcal{X}^{R_{i}})$;
// Impute missing data
for $tr\leftarrow 1$ to $T$ do
 for each $X^{R_{i}}_{(r)}\in\widetilde{\mathcal{X}}^{R_{i}}$ do
[…]
    // Domain transform
    Classify($\theta_{\text{Dis}}^{R_{i}}$, $X^{R_{i}\to S}_{(r)}$);
    if $tr=1\land r=1$ then
[…]
       Send($s$, $E^{R_{i}\to S}_{(r)}$);
    else
       if $tr=1\land r=2$ then
          $\overline{E}^{\mathcal{R}\to S}_{(r-1)}\leftarrow\textsc{Get}(s,r)$;
[…]
       else
          $\overline{X}^{\mathcal{R}\to S}_{(r-1)}\leftarrow\textsc{Get}(s,r)$;
       Classify($\theta_{\text{Dis}}^{R_{i}}$, $\overline{X}^{\mathcal{R}\to S}_{(r-1)}$);
       $X^{(\mathcal{R}\to S,R_{i})}_{(r)}\leftarrow\overline{X}^{\mathcal{R}\to S}_{(r-1)}+X^{R_{i}\to S}_{(r)}-X^{R_{i}\to S}_{(r-1)}$;
       // Masking
       Send($s$, $X^{(\mathcal{R}\to S,R_{i})}_{(r)}$);
Algorithm 1 The training of the FedTT framework in the client $c_{i}$
Input: Clients $\mathcal{C}=\{c_{1},\dots,c_{n}\}$ with source cities $\mathcal{R}=\{R_{1},\dots,R_{n}\}$; models $\theta_{\text{Dis}},\theta_{\text{TP}}$; local data $\mathcal{X}^{S}$
for $tr\leftarrow 1$ to $T$ do
 for $r\leftarrow 1$ to $R$ do
    if $tr=1\land r=1$ then
       $\{E^{R_{1}\to S}_{(r)},E^{R_{2}\to S}_{(r)},\dots\}\leftarrow\textsc{Get}(\mathcal{C},r)$;
       // Receive encrypted data
       $\overline{E}^{\mathcal{R}\to S}_{(r)}\leftarrow\sum_{i=1}^{n}E^{R_{i}\to S}_{(r)}$;
       // Aggregate encrypted data
[…]
    else
       $\{X^{(\mathcal{R}\to S,R_{1})}_{(r)},X^{(\mathcal{R}\to S,R_{2})}_{(r)},\dots\}\leftarrow\textsc{Get}(\mathcal{C},r)$;
       // Receive masked data
       $\overline{X}^{\mathcal{R}\to S}_{(r)}\leftarrow\sum_{i=1}^{n}X^{(\mathcal{R}\to S,R_{i})}_{(r)}-(n-1)\cdot\overline{X}^{\mathcal{R}\to S}_{(r-1)}$;
       // Aggregate masked data
[…]
 Classify($\theta_{\text{Dis}},\mathcal{X}^{S}$);
 // Update on server local data
 Prediction($\theta_{\text{TP}},\overline{\mathcal{X}}^{\mathcal{R}\to S},\mathcal{X}^{S}$);
 // Traffic prediction
Algorithm 2 The training of the FedTT framework in the server $s$